### v1

# Corpus Poisoning via Approximate Greedy Gradient Descent

Abstract
Dense retrievers are widely used in information retrieval and have also been successfully extended to other knowledge intensive areas such as language models, e.g., Retrieval-Augmented Generation (RAG) systems. Unfortunately, they have recently been shown to be vulnerable to corpus poisoning attacks in which a malicious user injects a small fraction of adversarial passages into the retrieval corpus to trick the system into returning these passages among the top-ranked results for a broad set of user queries. Further study is needed to understand the extent to which these attacks could limit the deployment of dense retrievers in real-world applications. In this work, we propose Approximate Greedy Gradient Descent (AGGD), a new attack on dense retrieval systems based on the widely used HotFlip method for efficiently generating adversarial passages. We demonstrate that AGGD can select a higher quality set of token-level perturbations than HotFlip by replacing its random token sampling with a more structured search. Experimentally, we show that our method achieves a high attack success rate on several datasets and using several retrievers, and can generalize to unseen queries and new domains. Notably, our method is extremely effective in attacking the ANCE retrieval model, achieving attack success rates that are 17.6% and 13.37% higher on the NQ and MS MARCO datasets, respectively, compared to HotFlip. Additionally, we demonstrate AGGD’s potential to replace HotFlip in other adversarial attacks, such as knowledge poisoning of RAG systems. [^1] Code can be find in https://github.com/JinyanSu1/AGGD
1 Introduction
Dense retrievers, despite their wide application and extensive deployment in real-world systems (Wan et al., 2022; Mitra et al., 2017; Lewis et al., 2020; Guu et al., 2020; Qu et al., 2021), have recently been shown to be vulnerable to various adversarial attacks such as corpus poisoning attacks (Zhong et al., 2023) and data poison attacks (Long et al., 2024; Liu et al., 2023), raising concerns about their security. Given that the corpora used in retrieval systems are often sourced from openly accessible platforms like Wikipedia and Reddit, a concerning scenario arises in which malicious actors can poison the retrieval corpus by injecting some adversarial passages, fooling the system into retrieving these malicious documents rather than the most relevant ones. Such attacks might be used for search engine optimization (Patil Swati et al., 2013) for promoting advertisement, or disseminating disinformation and hate speech.
A conventional approach for such attacks is HotFlip (Ebrahimi et al., 2018), which involves collecting a candidate set for a single randomly sampled token position and finding the best token in the candidate set with which to replace. In addition to corpus poisoning attacks on dense retrieval systems, HotFlip has been widely used in many other settings, such as knowledge poisoning attacks on retrieval augmented generation (RAG) systems (Zou et al., 2024) and adversarial prompt generation (Zou et al., 2023).
In this work, we begin by thoroughly investigating the HotFlip attack on dense retrieval systems to identify its limitations. Based on these insights, we propose a new general attack method called Approximate Greedy Gradient Descent (AGGD). Our experimental results show that AGGD can perform corpus poisoning attacks on dense retrieval systems more effectively, revealing their vulnerability. Though we use corpus attacks on dense retrievers as our primary example, it is important to note that AGGD can replace Hot Flip as a whole in any attack scenarios where HotFlip is applicable.
The main difference between AGGD and Hot Flip is that AGGD uses gradient information more effectively by selecting the top-ranked token from all token positions, rather than over a single randomly sampled position. This approach makes AGGD’s search trajectory deterministic, enabling a more structured best-first search. Experimental results demonstrate that AGGD achieves a high attack success rate across various datasets and retrieval models. In summary, our contributions are
  • We provide a thorough understanding of the existing HotFlip adversarial attack method, explaining its mechanics and identifying its potential problems.
  • We propose AGGD, a gradient-based method that replaces a randomized greedy search with a systematic best-first greedy search over the discrete token space. We demonstrate the effectiveness of AGGD in various settings.
  • We conduct extensive experiments to show the vulnerability of dense retrievers under AGGD. For example, when attacking the ANCE retriever, injecting just one adversarial passage (0.00003% of the NQ dataset corpus and 0.00001% of the MS MARCO dataset corpus) can achieve an attack success rate of 44.35% and 26.16% on these datasets, respectively, improving by 17.6% and 13.37% over HotFlip. The generated adversarial passage also possesses the capability to transfer to unseen queries in other domains, with AGGD achieving an attack success rate of 82.28% on the NFCorpus dataset.
2 Related Work
Dense Retrieval Dense retrievers utilize dense vector representations to capture the semantic information of passages and have demonstrated tremendous effectiveness compared to traditional retrieval systems (Yates et al., 2021). Consequently, they have been employed in many knowledge-intensive areas such as information retrieval (Karpukhin et al., 2020; Gillick et al., 2019; Wu et al., 2020; Wan et al., 2022; Mitra et al., 2017), open-domain question answering and language model pre-training (Lewis et al., 2020; Guu et al., 2020; Qu et al., 2021). For instance, retrieval-augmented generation (RAG) models (Lewis et al., 2020; Guu et al., 2020; Lee et al., 2019) combine language models with a retriever component to generate more diverse, factual and specific content.
Adversarial Attacks in Retrieval Systems Black-hat search engine optimization, which aims to increase the exposure of certain documents through malicious manipulation, poses a threat by reducing the quality of search results and inundating users with irrelevant pages (Castillo et al., 2011; Liu et al., 2023). Previous work has shown that retrieval systems are susceptible to small perturbations: making small edits to a target passage can significantly alter its retrieval rank (Song et al., 2020; Raval & Verma, 2020; Song et al., 2022) for individual or a small set of queries. More recently, a stronger setting known as corpus poisoning attack has been studied in Zhong et al. (2023), where the attack success rate of an adversarial passage is evaluated on unseen queries rather than on targeted given queries. These attacks differ from data poisoning attacks (Long et al., 2024; Liu et al., 2023; Chen et al., 2017; Schuster et al., 2020), as adversarial passages are injected into the retrieval corpus rather than the training data of the retrievers. The retrieval model remains unchanged in a corpus poisoning attack.
Discrete Optimization Many adversarial attacks in NLP involve discrete optimization, whether in classification tasks (Wallace et al., 2019; Ebrahimi et al., 2018; Song et al., 2021), retrieval systems (Jia & Liang, 2017; Song et al., 2020; Raval & Verma, 2020; Song et al., 2022), or adversarial prompt generation (Zou et al., 2023; Shin et al., 2020; Wen et al., 2024). The objective is to find small perturbations to the input to lead the model to make erroneous predictions. Due to the discrete nature of texts, directly applying adversarial attack methods from prior computer vision research (Xiao & Wang, 2021; Tolias et al., 2019) is infeasible. Instead, many methods build upon HotFlip (Ebrahimi et al., 2018) and approximate the effect of replacing a token using gradients.
3 Motivation
In this section we motivate our approach by formalizing the corpus poisoning problem setting (Section 3.1), describing and analyzing the standard HotFlip approach for producing adversarial passages (Zhong et al., 2023) (Section 3.2), and identifying a potential problem with this approach (Section 3.3).
3.1 Corpus Poisoning Problem Setting
In retrieval systems, the retrieval model takes a user query $q$ and returns a ranked list of the $k$ most relevant passages from a large corpus collection $\mathcal{C}=\{p_{1},\cdots,p_{|\mathcal{C}|}\}$ consisting of $|\mathcal{C}|$ passages. Compared to sparse retrieval models, which rely on lexical matching, dense retrievers rely on semantic matching. Specifically, the queries and the passages are first represented by $d$-dimensional dense vectors using a query encoder $E_{q}(\cdot)$ and passage encoder $E_{p}(\cdot)$, respectively. Relevance scores can then be computed according to a similarity function. A commonly used similarity function is the dot product of the dense vector representations of the query $q$ and the passage $p$: $\text{Sim}(q,p)=E_{q}(q)^{T}E_{p}(p)$. Finally, a ranked list of the $k$ most relevant passages $L=[\tilde{d}_{1},\tilde{d}_{2},\cdots,\tilde{d}_{k}],(L\subseteq\mathcal{C})$ is returned according to the relevance score.
We consider the problem of corpus attacks on a dense retrieval system, where we design an algorithm to find a small set of adversarial passages $\mathcal{A}=\{a_{1},\cdots,a_{|\mathcal{A}|}\}$ that can be retrieved by as many queries as possible for query distribution $\mathcal{P}_{\mathcal{Q}}$. These adversarial passages are then inserted into the corpus $\mathcal{C}$ to fool the dense retrieval models into retrieving passages from $\mathcal{A}$ rather than the most semantically relevant passages from the original corpus $|\mathcal{C}|$. The adversarial passage set $\mathcal{A}$ should be much smaller than the original corpus $\mathcal{C}$. The attack quality of the adversarial passage set $\mathcal{A}$ is typically evaluated based on its attack success rate, i.e., the percentage of queries for which at least one adversarial passage appears in the top-$k$ retrieval results.
Formally, the overall objective is to find an adversarial passage $a$, that maximizes the expected similarity to a query $q$ sampled from distribution $\mathcal{P}_{\mathcal{Q}}$, i.e.,
$a=\arg\max_{a}\mathbb{E}_{q\sim\mathcal{P}_{\mathcal{Q}}}\text{Sim}(q,a)$
In practice, we estimate the query distribution using a training set of queries $\mathcal{Q}=\{q_{1}.\cdots,q_{|\mathcal{Q}|}\}$, and we aim to find an $a$ with maximal similarity to $\mathcal{Q}$, i.e.,
$a=\arg\max_{a}\frac{1}{|\mathcal{Q}|}\sum_{q_{i}\in\mathcal{Q}}\text{Sim}(q_{i},a)=\arg\min_{a}\ell(a)$ (1)
where $\ell(a)=-\frac{1}{|\mathcal{Q}|}\sum_{q_{i}\in\mathcal{Q}}\text{Sim}(q_{i},a)$. The problem setting is realistic in search engines where a malicious user might perform search engine optimization to promote misinformation or spread spam.
Finding the exact solution to the optimization problem (1) is challenging since we are optimizing over a discrete set of inputs (i.e., the tokens in a passage). Additionally, running gradient descent on the embedding space might yield solutions that exist only in the embedding space and deviate significantly from valid texts in the discrete token space. In practice, a straightforward approach that leverages the gradient w.r.t. the one-hot token indicators can be employed to identify a set of promising candidates for replacement (Zou et al., 2023; Ebrahimi et al., 2018; Shin et al., 2020; Zhong et al., 2023). Specifically, we can compute the linearized approximation of replacing the $i$-th token $t_{i}$ in a passage $a$, by evaluating the gradient $\nabla_{e_{t_{i}}}\ell(a)$, where $e_{t_{i}}$ denotes the embedding of the token $t_{i}$. (Recall that sentence embeddings can be written as function of token embeddings, allowing us to compute the gradient with respect to the token embedding.) This idea has been adopted in many gradient-based search algorithms such as HotFlip (Ebrahimi et al., 2018) for producing adversarial texts, AutoPrompt (Shin et al., 2020) and Greedy Coordinate Gradient (GCG) (Zou et al., 2023) for generating prompts. We revisit the idea of HotFlip in the context of corpus attacks as an example.
3.2 HotFlip Revisited
In many text-based adversarial attacks, the goal is to find a perturbation of an input sequence — a randomly selected passage $a$ — to optimize some objective function. Due to the discrete nature of text, this is essentially a combinatorial search problem: the attacker searches over a class of perturbations on $a$ such as word swapping or character substitution (Morris et al., 2020) to produce the final adversarial text. The corpus poisoning approach of Zhong et al. (2023), for example, aims to fool a dense retrieval system to return adversarial passages for a broad set of user queries by transforming corpus passages (i.e., sequences of tokens) into adversarial passages $a$ that exhibit maximal similarity to a query set associated with the corpus.
Unfortunately, even if we fix the token length of a passage $|a|$ to be $m$, for a language model with vocabulary size $|\mathcal{V}|$, the total size of the search space is $|\mathcal{V}|^{m}$. For instance, if $m=30$, and $|\mathcal{V}|=32,500$, then the total number of sequences to evaluate is about $10^{135}$, which is computationally infeasible.
Thus, we need a method to identify a subset of the most promising perturbation sequences. The corpus poisoning attack by Zhong et al. (2023) addresses this problem using the HotFlip approach in a greedy search over candidate token perturbations. HotFlip (Ebrahimi et al., 2018) first randomly sample a token position, and then determines a candidate set of promising token replacements based on the gradient of their respective one-hot input vector [^2] The embedded version of the token is used. representations. Each iteration of search involves: (1) selecting a random position in $a$; (2) use gradient information to determine the top-$k$ token perturbation candidates; and (3) applying the perturbation (a token swap) that increases the similarity (decreases the loss) the most. [^3] If no candidate increases the loss, no change to $a$ is made.. This process continues for a fixed number of iterations.
3.3 Drawbacks of HotFlip
Although HotFlip usually works well in practice, there are some noticeable problems with this approach. For instance, it is less efficient due to its reliance on randomness for the search. Moreover, it is possible to try all token positions without finding any updates, causing HotFlip to get stuck and repeatedly search the same perturbation candidates. For example, consider an adversarial text with only two tokens (see Figure 2). If we first sample token position $t_{1}$ and observe no improvement for all candidates, there is still a $1/2$ probability of sampling the same token $t_{1}$ in the next iteration. Furthermore, if both token $t_{1}$ and token $t_{2}$ have been searched and show no improvement, the process would be stuck in a loop.
Figure 1: Comparing the true rank of words swapped with their rank according to the gradient-based Taylor approximation. The gradient identifies the top-1 correct token to swap 9% of the time, and guesses within the top ten tokens 58% of the time.
Figure 2: A simple example of finding a 2-token sequence $a$ through HotFlip (left) and AGGD (right). If HotFlip can’t find a better replacement for the currently sampled token position, there is a $\frac{1}{2}$ probability that it will sample the same token position again and redo the same evaluation, which is inefficient. Moreover, if the potential replacements for another token also don’t contain a better option, HotFlip continues to loop through the same search without reducing the loss.
Motivated by these existing problems, we propose a new algorithm called Approximate Greedy Gradient Descent (AGGD) that uses a deterministic greedy search that makes better use of gradient information by utilizing lower-ranked tokens(i.e., the overall most promising token swap candidates) to improve the quality of the candidate set. As shown in Figure 1, most high quality potential candidates are concentrated in the low-rank area. Suppose we aim to find a text with a token length $m=30$ by maintaining a candidate set of size $n=150$. Selecting as the candidate set the top-$5$ ranked gradients across all token positions is likely to result in better quality than selecting the top-$150$ ranked gradient candidates for only one token position. In the next section, we formally introduce our algorithm Approximate Greedy Gradient Descent (AGGD).
4 Approximate Greedy Gradient Descent
Similar to other gradient-based search algorithms, we fist initialize the adversarial passage to be $a=[t_{1},\cdots,t_{m}]$. We then iteratively update $a$ based on the best candidate that maximizes the similarity over batches of queries. Formally, at each step, we compute a first-order approximation of the change in the loss when swapping the $i$th token in $a$ with another token $v\in\mathcal{V}$. In contrast to Hotflip, AGGD identifies a candidate set for each of the $m$ tokens in $a$: for each token $i$, the candidate set contains the top $k=\frac{n}{m}$ ranked tokens from $\mathcal{V}$ according to the scoring function $s(v)=-e_{v}^{T}\nabla\ell(a)$, i.e., $\mathcal{X}_{i}=\text{top-}k_{v\in\mathcal{V}}[-e_{v}^{T}\nabla\ell(a)]$, where $e_{v}$ is the token embedding and the gradient is taken over the the embedding of the current adversarial passage $a$. Combining all the $\mathcal{X}_{i}$ leads to our overall candidate set $D^{(j)}=\cup_{i\in[m]}\mathcal{X}_{i}$ (as illustrated in Figure 3). For each candidate in the set $D$, the loss is re-evaluated and $a$ is updated to the candidate with lowest loss for the next step. This requires $n$ passes of the model, which constitutes the primary computational effort. If $a$ doesn’t update at iteration $i$, i.e., no candidate in $D^{(j)}$ achieves a lower loss, then in the next iteration, instead of searching over the top $k=\frac{n}{m}$, we search over the second tier of candidates, i.e., tokens from $\mathcal{V}$ with scores between top $\frac{n}{m}$ and the top $\frac{2n}{m}$, since the top $\frac{n}{m}$ candidates were already evaluated in the previous iterations. The search proceeds methodically as described above until a better candidate is found and $a$ is updated. The whole process is described in Algorithm 1.
Figure 3: Illustration of HotFlip (top) and AGGD (bottom) and their candidate sets.
Algorithm 1 Approximate Greedy Gradient Descent (AGGD)
  Input: The token length $m$ of the adversarial passage; Initialized adversarial passage $a=[t_{1},\cdots,t_{m}]$; Total number of iterations $N$; Size of the candidate set $n$; Search depth $d=0$.
  Let $k=\frac{n}{m}$ to be the per token candidate set size.
  for $j=0,1,2,\cdots,N$  do
     for $i=1,\cdots,m$ do
        Let $\mathcal{V}_{o}=\mathcal{R}(s(a))$ where $s(a)=-e_{v}^{T}\nabla\ell(a)$, $\ell(a)=-\frac{1}{|\mathcal{Q}|}\sum_{q}\text{Sim}(q,a)$ and $\mathcal{R}(\cdot)$ is a rank function.
        Let $\mathcal{X}_{i}=\mathcal{V}_{o}[(d-1)k,dk]$ be the truncated top-$(d-1)k$ to top-$dk$ candidates.
     end for
     Construct candidate set at iteration $j$ as $D^{(j)}=\cup_{i\in[m]}\mathcal{X}_{i}$
     Let $a^{{}^{\prime}}=\arg\min_{a^{{}^{\prime}}\in D^{(j)}}\ell(a^{{}^{\prime}})$ // Check if a better adversarial passage exists in candidate set
     if $\ell(a^{{}^{\prime}})<\ell(a)$ then
        Update $a=a^{{}^{\prime}}$, reset depth $d=0$. // Update the adversarial passage if find a better one
     else
        Update $d=d+1$ // If no update, search from top-$(d-1)k$ to top-$dk$ in next iteration
     end if
  end for
  Return Final adversarial passage $a$
5 Experiments
5.1 Experimental Details
Datasets We primarily use two popular question-answering datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019) and MS MARCO (Nguyen et al., 2016) for our attack. NQ, containing 132,803 question-answer pairs, is collected from Wikipedia, while MS MARCO, containing 532,761 question answer pairs, is collected from web documents. For in-domain attacks, we evaluate the attack on the unseen queries on held-out test queries of these two datasets. To test the transferability of our attack, we also evaluate it on 5 out-of-domain datasets: NFCorpus Boteva et al. (2016), Quora, SCIDOCS Cohan et al. (2020), SciFact Wadden et al. (2020), FiQA-2018 Maia et al. (2018). These datasets contain unseen queries and corpora that are out of distribution or from entirely different domains such as biomedicine, scientific articles, and finance. Statistics of these datasets can be found in Appendix C.1.
Retrievers In our main experiments, we conduct attacks on 5 state-of-the-art retrieval models: Contriever, Contriever-MS (Contriever fine-tuned on MS MARCO) Gautier et al. (2022), DPR-nq (trained on NQ), DPR-mul (trained on multiple datasets) Karpukhin et al. (2020) and ANCE Xiong et al. (2020).
Evaluation Metrics After generating the adversarial passages on the training set and injecting them into the corpus, we evaluate the effectiveness of our attack using the top-$k$ attack success rate (ASR) on test queries. Top-$k$ ASR is defined as the percentage of queries for which at least one adversarial passage is retrieved in the top-$k_{r}$ results, i.e., $\text{ASR}=\frac{1}{n_{q}}\sum_{i=1}^{n_{q}}\mathbbm{1}\{a\in\mathcal{R}_{r}(q^{\text{test}}_{i},k_{r})\}$, where $n_{q}$ is the total number of test queries and $\mathcal{R}_{r}(q_{i}^{\text{test}},k_{r})$ is the retriever that returns the top-$k_{r}$ most relevant passages for the test test query $q_{i}^{\text{test}}$. $\mathbbm{1}\{a\in r(q_{i}^{\text{test}},k_{r})\}$ is the indicator function, which equals to 1 if $a\in\mathcal{R}(q_{i}^{\text{test}},k_{r})$ and 0 otherwise. A higher ASR indicates that the model is more vulnerable to attacks, and thus, the attack is more effective. We use $k_{r}=1000$ unless specified otherwise.
Additionally, we use retrieval accuracy on the validation data, which is defined as
$\text{RetAcc}=\frac{1}{n_{\text{val}}}\sum_{i=1}^{n_{\text{val}}}\mathbbm{1}\{\text{Sim}(q_{i}^{\text{val}},p_{i}^{\text{val}})>\text{Sim}(q_{i}^{\text{val}},a)\}$
where $a$ is the adversarial passage, $n_{\text{val}}$ is the total number of query-passage pairs in the validation set, and $p_{i}^{\text{val}}$ is most semantically relevant passage that should have been retrieved by query $p_{i}^{\text{val}}$. Lower retrieval accuracy indicates a higher chance of the model being fooled into choosing the adversarial texts $a$ over the most semantically relevant passage in the corpus. Note that the Success Rate during training (on validation data) can be defined through retrieval accuracy, which is simply $1-\text{RetAcc}$.
Hyperparameters For all baselines, we use adversarial passages of $m=30$ tokens and perform the token replacement for 2000 steps. We fix the candidate set size at $n=150$. All the experiments are conducted on NVIDIA A100 GPU (with total memory 40G).
5.2 Main Results
Dataset Method Retriever
Contriever Contriever-MS ANCE DPR-mul DPR-nq
NQ AGGD 89.57(3.58) 54.53(11.29) 44.35(3.54) 1.47(0.41) 0.14(0.06)
HotFlip 88.46(0.84) 49.64(6.09) 26.75(1.45) 1.01(0.15) 0.19(0.12)
Random 74.64(1.78) 18.34(1.48) 7.75(1.88) 0.5(0.21) 0.19(0.1)
MS MARCO AGGD 75.58(5.81) 13.95(15.16) 26.16(13.3) 2.91(1.01) 2.91(1.93)
HotFlip 75.0(4.76) 13.95(9.16) 12.79(5.07) 1.74(1.01) 0.58(1.01)
Random 55.81(6.98) 1.74(3.02) 5.81(3.49) 1.16(2.01) 0.58(1.01)
Table 1: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO datasets with 5 retrievers by injecting 1 adversarial passage. (Equivalent to increasing the corpus size by 0.00003% for NQ and 0.00001% for MS MARCO.) We highlight the best performing attacking method in bold (Higher ASR indicates better attack performance). Results are from 4 random runs with standard deviation in parenthesis.
In-Domain Attack We show the in-domain attack performance of AGGD and the other two baselines (HotFlip and random perturbation) in Table 1, evaluating the injection of only one adversarial passage. Our findings are as follows: (1) The pretrained Contriever model is more vulnerable to attacks. All three attack baselines achieve the highest ASR compared to other retrieval models. Poisoning with AGGD achieves 89.57% ASR on NQ dataset and 75.58% on MS MARCO dataset, respectively. Even using Random perturbation can achieve a relatively high ASR of 74.64% on NQ dataset and 55.81% on MS MARCO dataset. (2) Besides achieving comparable results in Contriever and DPR, AGGD is extremely effective in attacking ANCE, improving over the second-best baseline by 17.6% and 13.37% on NQ and MS MARCO, respectively. The effectiveness of AGGD in attack ANCE can also be clearly observed during the training. Due space limitations, we provide training accuracy in Appendix A.2(Figure 10). (3)Supervised retriever models such as DPR are more challenging to attack. Attacking DPR-nq on NQ with only 1 adversarial passage does no better than random perturbation. (4) MS MARCO seems to be harder to attack than NQ, likely due to its larger testing corpus, and thus, lower poison rate.
Target Domain Source Domain Methods Retriever
Contriever Contriever-MS ANCE DPR-mul DPR-nq
FiQA-2018 NQ AGGD 62.04(17.22) 3.78(1.39) 4.09(1.18) - -
HotFlip 65.28(9.06) 2.08(0.55) 2.89(0.62) 0.23(0.23) 0.15(0.11)
Random 37.38(2.76) 0.27(0.13) 0.96(0.33) 0.15(0.11) 0.31(0.11)
MS MARCO AGGD 84.57(1.97) 14.16(21.18) 11.46(1.61) 1.89(0.56) 1.97(1.24)
HotFlip 75.5(7.87) 3.24(2.06) 7.02(0.56) 1.2(0.6) 1.81(1.18)
Random 58.76(6.12) 0.23(0.32) 3.51(1.06) 0.46(0.24) 0.42(0.13)
NFCorpus NQ AGGD 51.32(7.79) 20.98(5.59) 49.54(6.21) 12.62(2.09) 16.25(0.56)
HotFlip 45.9(11.47) 13.24(2.6) 46.13(2.22) 13.85(3.57) 17.41(1.63)
Random 28.79(6.24) 8.67(2.19) 27.01(5.86) 10.29(2.03) 16.18(0.46)
MS MARCO AGGD 56.58(14.85) 22.91(18.3) 82.28(3.41) 38.16(5.66) 40.63(4.49)
HotFlip 54.49(6.28) 17.96(10.63) 65.56(4.99) 29.18(3.16) 34.98(8.55)
Random 32.97(5.71) 8.59(8.34) 55.73(7.11) 23.61(4.15) 23.14(2.81)
Quora NQ AGGD 57.36(14.93) 11.24(4.69) 24.48(3.48) 0.36(0.06) 0.27(0.14)
HotFlip 61.82(4.97) 8.09(4.42) 20.43(0.73) 0.3(0.14) 0.2(0.06)
Random 30.09(4.01) 1.03(0.1) 10.88(0.94) 0.17(0.02) 0.27(0.05)
MS MARCO AGGD 73.42(4.37) 8.25(12.74) 37.07(2.52) 3.32(1.32) 3.37(1.07)
HotFlip 69.62(8.69) 2.75(1.59) 27.92(2.31) 2.03(0.95) 2.31(0.34)
Random 42.14(7.59) 0.37(0.45) 18.39(0.85) 0.8(0.2) 0.7(0.54)
SCIDOCS NQ AGGD 21.18(7.91) 11.75(7.84) 9.55(1.72) 0.87(0.26) 0.4(0.1)
HotFlip 27.15(13.45) 14.1(4.41) 7.18(1.0) 1.18(0.4) 0.35(0.11)
Random 12.82(3.62) 0.73(0.58) 2.9(0.8) 0.57(0.37) 0.27(0.08)
MS MARCO AGGD 23.25(6.88) 14.22(20.72) 31.32(7.33) 3.92(1.7) 1.13(0.45)
HotFlip 24.98(6.23) 11.98(7.4) 21.25(4.37) 3.05(1.34) 0.85(0.32)
Random 15.02(1.6) 1.23(1.74) 10.58(5.27) 2.35(1.27) 0.23(0.08)
SciFact NQ AGGD 25.83(4.19) 1.58(0.86) 10.83(2.88) 0.17(0.17) -
HotFlip 28.83(14.18) 1.17(0.83) 6.75(0.92) 0.42(0.43) -
Random 9.58(4.44) - 1.42(0.86) - -
MS MARCO AGGD 31.83(10.62) 3.75(5.01) 48.0(7.04) 3.5(0.9) 1.58(0.49)
HotFlip 25.0(7.06) 2.5(1.89) 28.17(8.07) 2.83(0.5) 1.42(1.44)
Random 12.08(6.19) 0.17(0.17) 13.42(7.98) 1.75(1.09) 0.67(0.71)
Table 2: Out-of-domain top-50 attack success rate with only 1 adversarial passage. Results are averaged over 4 random runs with standard deviation shown in parenthesis. The combinations of attack and source dataset that achieve the highest ASR on each target domain and retriever are highlighted in bold.
Out-of-Domain Attack We found that the generated adversarial passages can transfer across different domains. In Table 2, we use adversarial passages generated from training set of NQ and MS MARCO and insert them into the corpora of retrieval tasks in other domains. Since the out-of-domain datasets have smaller corpora compared to NQ and MS MARCO, we use top $k_{r}=50$ ASR (rather than $k_{r}=1000$ used for in-domain attacks) to measure the effectiveness of the attack method. We found that (1) Compared to adversarial passages generated from NQ dataset, those trained from MS MARCO generally perform better in out-of-domain attacks, possibly because MS MARCO contains more training data. (2) Contriever models are still the most vulnerable to corpus poisoning attacks. For example, inserting a single adversarial passage generated by AGGD into FiQA-2018 achieves a top-$k_{r}=50$ ASR of 84.57%, and inserting it into Quora can trick the model into returning the adversarial passage in the top-50 retrieved passage for 73.42% of the queries. (3) NFCorpus is an easier domain to attack. AGGD achieves over 38% ASR even in DPR-mul and DPR-nq, which is surprising since, as shown in the in-domain attack results in Table 1, attacking DPR-mul and DPR-nq is extremely hard , even with adversarial passage trained from in-domain data.
5.3 Analysis and Ablation Study
(a) Attack Success rate of candidates sets collected by different methods. (Averaged over 400 candidate sets sampled.)
(b) The proportion of times the best candidate occurs in the candidate sets collected by AGGD, HotFlip, and Random.
Figure 4: Experiments on Contriever with NQ dataset illustrate that the candidate set collected by AGGD has higher overall quality (left) and is more likely to contain the best candidate (right).
Candidate Set Quality Our results indicate that candidate sets selected using AGGD have better quality than those selected using HotFlip and Random perturbation. To demonstrate this, we randomly sampled a passage and used AGGD, HotFlip and random perturbation to select their respective candidate sets with $m=150$ candidates. We then evaluated the success rate (i.e., $1-\text{RetAcc}$) on the validation set for all the candidates. The results, averaged from 400 random samples, are shown in Figure 4(a). We can observe that the candidate set from AGGD has higher mean success rate on validation data. Additionally, the higher confidence bound in the AGGD candidate set signifies that AGGD’s candidate set not only has higher overall quality, but is also more likely to contain the best candidate when compared horizontally across candidate sets from other methods. To verify this, we counted how often the best candidate occurs in candidate sets selected by these methods and found that more than 85% of the time, the AGGD candidate set contains the best candidate, while less than $13\%$ show up in HotFlip candidate set, as shown in Figure 4(b). More experiments on other retrievers models further verify that the AGGD candidate set has higher quality (Figure 13 in Appendix B.1).
The Effects of Candidate Set Size $n$
Figure 5: The effect of candidate set size $n$ on the attack success rate.
In Figure 5, we show how the candidate set size $n$ affects the ASR evaluated on NQ dataset with three retrievers. Generally, we can observe that increasing size of the candidate set improves the ASR. The effect of the candidate set size would be more pronounced when a larger range of candidate set sizes is considered.
The effect of the token length $m$
Figure 6: The effect of the number of tokens, with fixed candidate set size $n=100$ and varying adversarial passage token length $m=\{5,10,25,50,100\}$.
In Figure 6, we show the attack success rate for varying token lengths $m=\{5,10,25,50,100\}$ while fixing the candidate set size $n=100$. Generally, we find that larger token lengths result in higher attack success rate upon convergence. This is intuitive since larger $m$ indicates a larger subspace in the continuous dense embedding space, thus providing a higher lower bound on the effectiveness of the adversarial passage we can find.
5.4 Extending AGGD to Knowledge Poisoning Attacks
Similar to Hot Flip, which can be used for many other adversarial attacks such as adversarial prompt generation Zou et al. (2023) and knowledge poisoning in RAG Zou et al. (2024). AGGD can also be conveniently used as a plug-in replacement for HotFlip in these attacks. In Table 3, we experiment with PoisonedRAG Zou et al. (2024) in their white box setting, where Hot Flip was originally used to craft adversarial texts. The details in reproducing Table 3 are given in Appendix D. We find that AGGD achieves a comparable or better attack success rate than Hot Flip in both MS MARCO and NQ dataset on multiple LLMs.
LLaMa-2-7B LLaMa-2-13B Vicuna-7B Vicuna-13B Vicuna-33B GPT-3.5 GPT-4
MS MARCO
AGGD 0.81(0.00) 0.79(0.01) 0.73(0.01) 0.72(0.01) 0.67(0.01) 0.80(0.01) 0.88(0.01)
HotFlip 0.80(0.02) 0.77(0.02) 0.69(0.02) 0.73(0.04) 0.69(0.02) 0.79(0.01) 0.86(0.01)
NQ
AGGD 0.85(0.00) 0.94(0.00) 0.81(0.00) 0.82(0.00) 0.68(0.00) 0.92(0.00) 0.97(0.00)
HotFlip 0.83(0.00) 0.95(0.00) 0.80(0.00) 0.81(0.00) 0.67(0.00) 0.87(0.01) 0.96(0.01)
Table 3: Comparing Attack success rate of PoisonedRAG (white-box) using HotFlip and AGGD to find the adversarial texts.
6 Conclusion
In this paper, we propose a new adversarial attack method called AGGD, a gradient-based search algorithm that systematically and structurally finds potential perturbations to optimize the objective function. We use the corpus poisoning attacks as the main example to demonstrate the effectiveness of our algorithm. Experiments on multiple datasets and retrievers show that the proposed approach is effective in corpus poisoning attacks, achieving high attack success rate in both in-domain and out-of-domain scenarios, even with an extremely low poison rate. Additional experiments on other adversarial attacks indicate the potential of AGGD as a competitive alternative to the widely used HotFlip.
7 Limitations and Broader Impact
Though we use corpus poisoning attack as the main example to deliver our attack algorithm AGGD, our proposed attack can be used for various attack scenarios, thus, we didn’t specifically investigate defenses of AGGD for corpus poison attack. Since our method assembles to HotFlip, existing defenses to these attacks should be able to defend our attack to some extend. However, as our attack is stronger than HotFlip, it remains to understand how well these defense works.
Appendix A Additional Experiments
A.1 Comparing HotFlip with Random Perturbation
As illustrated in Figure 7, HotFlip first randomly selects a token position for replacement and identifies $n$ potential alternatives for that position. It updates the adversarial passage with the most effective perturbation from this candidate set if it improves attack performance. If no perturbation enhances performance, HotFlip progresses to the next iteration by randomly selecting another token position and repeating the process. It is useful then to compare Hotflip vs. a natural black box Random approach to candidate generation. As depicted in Figure 7), the Random approach creates a candidate set by randomly sampling from the vocabulary.
Figure 7: Comparing HotFlip attack and the random perturbation attack. The main difference is how they select the candidate set. HotFlip uses the gradient information to enhance the quality of the candidate set, whereas random perturbation selects candidates uniformly from $\mathcal{V}$ without leveraging gradient information.
HotFlip is more Efficient Than Random Perturbation
In experiments using three dense retrievers on the widely used Natural Questions (NQ) dataset, we find that HotFlip outperforms the Random perturbation baseline, as shown in Figure 8. Specifically, HotFlip consistently achieves low retrieval accuracy on the validation set while maintaining candidate sets of the same size or smaller. (Red lines are mostly lower than the blue lines.) Equivalently, to achieve the same degree of attack effect, HotFlip is more efficient because it requires a smaller candidate set. For example, as depicted in Figure 9, Random perturbation needs to maintain a candidate set of size $n=900$ to match the performance of HotFlip with only $n=30$ candidates. This means that the size of Random perturbation candidate set is 30 times larger and consequently, Random perturbation is 30 times slower.
HotFlip and Random perturbation share the same intuition of maintaining a candidate set and then take a greedy token swap based on this set. Though many factors can influence the efficiency and the effectiveness the greedy search, one of the major factors is the quality of the candidate set. If the candidate set is of high quality, we can maintain a smaller size candidate set, reducing the number of searches at each greedy step. For example, HotFlip filters down the most likely potential tokens from $|\mathcal{V}|$ to $n$ based on the gradient information, while random perturbation samples $n$ tokens from $\mathcal{V}$. As a result, random perturbation has a lower quality candidate set. Therefore, it is crucial to improve the quality of the candidate set.
Figure 8: Retrieval accuracy of HotFlip and random perturbation on the NQ dataset with varying candidate set size $n=\{30,60,90,150\}$. Lower retrieval accuracy suggests a more successful adversarial passage attack, as evaluated on the validation set during greedy search iterations.
Figure 9: Retrieval accuracy of HotFlip and random perturbation. Hot Flip only needs to maintain a candidate set of size $n=30$ to achieve the same performance of random perturbation with a candidate set size $n=900$.
A.2 Retrieval Accuracy during Training
Retrieval Accuracy for 1 Adversarial Passage
In Figure 10, we plot the Retrieval accuracy during training on both NQ and MS MARCO datasets for all 5 retrievers as a complement to the results in Table 1. During training, at each iteration, we evaluate the retrieval accuracy of the best adversarial passage $a_{i}$ from the candidate set, and compare it with the current adversarial passage to decide whether to update it. Due to the greedy nature of these attacks, Retrieval accuracy is guaranteed to descend. Figure 10 shows that, for both NQ and MS MARCO dataset, the harder-to-attack models such as ANCE, DPR-mul and DPR-nq converge with high retrieval accuracy. In contrast, for the easier-to-attack models such as Contriever, the attack methods have not yet converged, even though the retrieval accuracy is as low as around 0.2.
Figure 10: Retrieval accuracy (the portion of the queries that have higher similarity to the adversarial passage than the gold passage) on validation data during training. Lower retrieval accuracy indicates higher chance of adversarial passage being retrieved.
Effect of the Candidate Set Size to Retrieval Accuracy
Figure 11: The retrieval accuracy during training with various candidate set sizes. ($n=30,60,90,150)$.
In Figure 11, we illustrate the retrieval accuracy during training with varying candidate set sizes on NQ dataset and 3 retrievers. Larger candidate set sizes generally lead to lower retrieval accuracy.
Effect of Token Length on Retrieval Accuracy
In Figure 12, we show the training retrieval accuracy with various token length settings. We find that, within a proper range, longer token lengths lead to more effective adversarial passages.
Figure 12: The retrieval accuracy during training with various token lengths ($m\in\{5,10,25,50,100\}$) and a fixed candidate set size $n=100$.
Appendix B Experimental Results with 10 Adversarial Passages
To generate multiple adversarial passages, we follow Zhong et al. (2023) by clustering similar queries based on their embeddings using the $k$-means algorithm. We then generate one adversarial passage for each cluster.
B.1 Candidate Set Quality
In Figure 13, we present additional experiments with two more retriever models: Contriever-MS and ANCE. We observed a similar trend, over the 400 random samples, with more than 92% of the best candidate belonging to AGGD candidate set. This further supports the conclusions in Section 5.3.
(a) Attack Success rate of candidate sets collected by different methods. (Results are averaged over 400 candidate sets sampled when training with Contriever-MS (left) and ANCE (right) on NQ dataset.)
(b) The portion of the times the best candidate occurs in the candidate sets collected by AGGD, HotFlip and Random, respectively, when training on NQ dataset with Contriever-MS (left) and ANCE (right).
Figure 13: Additional experiments on Contriever-MS and ANCE with NQ dataset illustrate that candidate sets collected by AGGD (a) has higher overall quality and (b) are more likely to contain the best candidate.
In-Domain Attack with 10 Adversarial Passages
Table 4 shows the in-domain attack results of inserting 10 adversarial passages into the NQ and MS MARCO datasets, with a poison rate of around $0.0003\%$ of the NQ corpus and $0.0001\%$ of the MS MARCO corpus. Similar to injecting only 1 adversarial passage, we found that (1) The pretrained Contriever model is easy to attack: even with just 10 adversarial passages, all three baselines (AGGD, HotFlip and Random perturbation) successfully attack more than 90% of the queries in both NQ and MS MARCO datasets, tricking the Contriever into returning the adversarial passage among the top-1000 retrieved results. For NQ dataset, even random perturbation achieves a high ASR of 97.75%, while the other two methods using gradient information achieves higher ASR of $>99\%$. (2) AGGD still outperforms HotFlip on ANCE, with improvements of 11.95% and 25.58% on NQ and MS MARCO datasets, respectively. (3) Though attacking DPR-nq on NQ with only 1 adversarial passage does no better than random perturbation (as shown in Table 1), increasing the number of adversarial passage to 10 allows both AGGD and HotFlip to perform better than random perturbation, with ASR improvement of 10.08% and 6.67% over the Random perturbation baseline.
Dataset Methods Retriever
Contriever Contriever-MS ANCE DPR-mul DPR-nq
NQ AGGD 99.46(0.01) 86.76(1.48) 85.53(0.42) 23.41(0.03) 16.87(0.07)
HotFlip 99.06(0.07) 86.72(2.25) 73.58(1.97) 25.77(1.32) 13.46(0.42)
Random 97.75(0.04) 78.06(2.56) 55.87(1.35) 15.34(0.74) 6.79(0.07)
MS MARCO AGGD 96.51(1.16) 54.65(1.16) 72.09(0.0) 13.95(0.0) 5.81(3.49)
HotFlip 96.51(3.49) 59.3(3.49) 46.51(2.33) 11.63(0.0) 2.33(0.0)
Random 91.86(1.16) 27.91(2.33) 30.23(9.3) 5.81(3.49) 1.16(1.16)
Table 4: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO dataset with 5 retrievers by injecting 10 adversarial passages. (adding 0.0003% corpus to NQ and 0.0001% to MS MARCO dataset). The best-performing attacking method is highlighted with in bold (higher ASR indicates better attack performance).
Out-of-Domain Attack with 10 Adversarial Passages
In Table 5, we perform the out-of-domain attack by inserting more adversarial passages. We find that, inserting more adversarial passages significantly improves the attack transferability of HotFlip, enabling it to outperform AGGD in models such as Contriever-MS, DPR-mul and DPR-nq. However, AGGD still performs much better than HotFlip when using ANCE as the retriever.
Target Domain Source Domain Methods Retriever
Contriever Contriever-MS ANCE DPR-mul DPR-nq
FiQA-2018 NQ AGGD 90.97(0.08) 25.08(7.02) 12.27(0.39) 4.32(0.0) 2.7(0.85)
HotFlip 79.55(1.77) 20.14(2.86) 8.33(1.7) 1.77(1.31) 1.62(0.54)
Random 79.17(0.0) 14.43(4.55) 5.56(0.31) 1.23(0.0) 1.23(0.15)
MS MARCO AGGD 97.38(0.31) 45.99(25.31) 25.69(1.16) 21.22(7.95) 14.35(6.33)
HotFlip 94.06(1.47) 50.08(3.32) 18.75(0.39) 13.81(5.17) 20.52(0.77)
Random 83.41(0.23) 5.25(0.31) 8.33(0.93) 10.57(5.32) 4.71(3.01)
NFCorpus NQ AGGD 83.13(2.01) 43.5(2.32) 75.54(3.72) 25.23(6.97) 25.23(2.94)
HotFlip 88.85(0.62) 42.88(3.56) 63.0(1.39) 31.89(1.24) 27.09(0.16)
Random 68.27(0.46) 29.72(4.02) 48.3(2.48) 22.29(0.93) 21.21(2.01)
MS MARCO AGGD 97.06(0.15) 62.07(25.85) 95.2(0.46) 72.29(7.9) 64.24(4.49)
HotFlip 90.56(4.8) 78.02(9.6) 89.78(0.93) 69.35(11.15) 74.77(5.42)
Random 88.85(3.1) 63.78(1.86) 81.58(1.7) 47.21(3.87) 61.92(8.98)
Quora NQ AGGD 85.62(0.13) 35.84(2.36) 42.12(0.82) 2.7(0.32) 3.53(0.89)
HotFlip 77.47(1.67) 32.46(9.31) 33.01(2.55) 3.84(0.03) 4.04(0.13)
Random 67.66(5.7) 23.15(5.66) 26.36(0.42) 1.64(0.09) 2.36(0.47)
MS MARCO AGGD 91.8(2.98) 34.08(5.24) 48.4(0.52) 12.75(0.04) 12.0(1.43)
HotFlip 92.31(3.08) 40.83(3.2) 41.58(0.04) 12.93(0.48) 13.66(0.03)
Random 65.28(3.2) 11.79(3.52) 30.69(1.66) 6.43(0.13) 7.09(0.95)
SCIDOCS NQ AGGD 53.45(7.45) 45.7(17.7) 33.0(0.9) 1.75(0.35) 0.45(0.05)
HotFlip 49.7(2.2) 40.95(5.15) 25.4(5.7) 3.2(0.1) 1.0(0.2)
Random 37.7(4.1) 17.95(1.35) 13.25(0.95) 2.2(0.5) 0.55(0.25)
MS MARCO AGGD 65.4(2.5) 41.0(10.1) 45.25(2.15) 24.5(2.2) 6.4(0.0)
HotFlip 51.35(4.65) 67.55(8.25) 36.1(0.2) 21.85(2.95) 7.95(0.75)
Random 39.05(4.55) 13.6(4.1) 20.25(0.65) 14.7(2.9) 2.5(1.1)
SciFact NQ AGGD 66.83(5.5) 16.33(9.67) 56.83(9.5) 1.83(1.5) 0.33(0.33)
HotFlip 63.5(12.17) 15.67(3.0) 27.0(8.0) 2.83(1.17) 1.0(0.0)
Random 41.17(4.5) 5.0(3.0) 12.67(0.67) 1.0(0.0) 0.67(0.0)
MS MARCO AGGD 93.83(2.5) 49.83(13.17) 72.5(1.17) 17.33(5.67) 17.5(10.5)
HotFlip 89.5(3.17) 33.5(8.83) 58.83(6.17) 24.5(1.17) 21.17(10.83)
Random 80.83(8.17) 13.67(4.0) 32.67(8.0) 8.33(2.33) 14.5(2.17)
Table 5: Out-of-domain top-50 attack success rate with 10 adversarial passages. Due to the computational constraints, results are averaged over 2 random runs, with standard deviations shown in parentheses. The combinations of attack and source dataset that achieve the highest ASR for each target domain and retriever are highlighted in bold.
Appendix C Experimental Detail
Domain Dataset Train Dev Test
# Pair # Query # Query # Corpus
In-domain Web MS MARCO Nguyen et al. (2016) 532,761 - 6,980 8,841,823
Wikipedia NQ Kwiatkowski et al. (2019) 132,803 - 3,452 2,681,468
Bio Medical NFCorpus Boteva et al. (2016) 110,575 324 323 3,633
Out-of-domain Quora Quora - 5,000 10,000 522,931
Scientific SCIDOCS Cohan et al. (2020) - - 1,000 25,657
SciFact Wadden et al. (2020) 920 - 300 5,183
Finance FiQA-2018 Maia et al. (2018) 14,166 500 648 57,638
Table 6: Dataset Statistics. More statistics can be found in Zhao et al. (2024); Thakur et al. (2021). In our experiments, we use MS MARCO and NQ datasets to train the adversarial passages and evaluate the attack on the unseen test queries from these two datasets for in-domain attack evaluation. The remaining 5 datasets (NFCorpus, Quora, SCIDOCS, SciFact, FiQA-2018) are used for out-of-domain evaluation when injecting the adversarial passages generated from the MS MARCO and NQ dataset into the corpora of these out of domain datasets.
C.1 Dataset
  • MS MARCO Nguyen et al. (2016) contains a large amount of queries with annotated relevant passages from Web documents.
  • Natural Questions (NQ) Kwiatkowski et al. (2019) contains Google search queries with annotations from the top-ranked Wikipedia pages.
These two datasets have been widely used for evaluating dense retrieval models.
The statistics of all the datasets used in evaluation are summarized in Table 6.
C.2 Retrievers
We experimented with the following retrievers:
  • Dense Passage Retriever(DPR) Karpukhin et al. (2020) is a two-tower bi-encoder trained with a single BM25 hard negative and in-batch negatives. It has been used as the retrieval component of many Retrieval-Augmented Generation (RAG) models Lewis et al. (2020). In our paper, we use both the open-sourced Multi model (DPR-mul), which is a bert-base-uncased model trained on four QA datasets (NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions Berant et al. (2013) and CuratedTREC Baudiš & Šedivỳ (2015)) and the single NQ model (DPR-nq).
  • ANCE Xiong et al. (2020) is a bi-encoder that generates hard negatives using an approximate Nearest Neighbor (ANN) index of the corpus. The index is continuously updated in parallel to identify challenging negative examples for the model during fine-tuning.
  • Contriever Gautier et al. (2022) is an unsupervised dense retriever using contrastive learning. It leverages the BERT architecture to encode both queries and documents. Contriever-MS (Contriever fine-tuned on MS MARCO) is a version of the Contriever model that has been fine-tuned using the MS MARCO dataset, which provides large-scale, supervised training data.
Appendix D Knowledge Poisoning Attacks to RAG
LLaMa-2-7B LLaMa-2-13B Vicuna-7B Vicuna-13B Vicuna-33B GPT-3.5 GPT-4
MS MARCO
AGGD 0.92(0.00) 0.92(0.00) 0.92(0.00) 0.92(0.00) 0.92(0.00) 0.92(0.00) 0.92(0.00)
HotFlip 0.90(0.01) 0.90(0.01) 0.90(0.01) 0.90(0.01) 0.90(0.01) 0.90(0.01) 0.90(0.01)
NQ
AGGD 1.00(0.00) 1.00(0.00) 1.00(0.00) 1.00(0.00) 1.00(0.00) 1.00(0.00) 1.00(0.00)
hotflip 1.00(0.00) 1.00(0.00) 1.00(0.00) 1.00(0.00) 1.00(0.00) 1.00(0.00) 1.00(0.00)
Table 7: PoisonedRAG performance on multiple dataset and LLMs using Hot Flip and AGGD evaluated using F1 score.
Retrieval-Augmented Generation
Karpukhin et al. (2020); Lewis et al. (2020); Borgeaud et al. (2022); Thoppilan et al. (2022) augments LLMs with external knowledge retrieved from a knowledge base to improve their ability to generate accurate and up-to-date content. There are three components in RAG: knowledge base, the retriever and the LLM. The knowledge base contains a large corpus collected from various domains such as Wikipedia Thakur et al. (2021), Fiance Loukas et al. (2023) and Biomedical articles Roberts et al. (2020). Given a user question, the retriever uses a text encoder to compute the embedding vector. A set of $k$ retrieved texts from the knowledge base with the highest similarity to the question are then retrieved, which can be used by the LLM to generate content.
RAG enables LLMs to incorporate more current knowledge by regularly updating the knowledge base. However, this also introduces potential security concerns: maliciously crafted content could be injected into the database during updates, which might then be retrieved by the LLM to generate false, biased or harmful output.
Threat Model
We consider a scenario where the attacker does not have direct access to the database, but can access the weights of the retrieval model. This scenario is realistic, as the database is likely hosted on a secure system, while the retriever used may be an open-access LLM. We assume the attacker can inject a few carefully crafted corpus into the knowledge base.
Metric
We use the same metrics, Attack Success Rate (ASR) and F1, as used in Zou et al. (2024). However, since ASR is task-specific, it might not be exactly the same as defined for corpus poisoning attacks. In PoisonedRAG, the definitions are given as follows.
  • ASR The fraction of target questions whose answers are the attacker-chosen target answers.
Additionally, we report average F1-Score over different target questions in Table 7, which measures the trade-off between Precision and Recall. Specifically,
[…]
where Precision is defined as the fraction of poisoned texts among the top-$k$ retrieved ones for the target question and Recall is defined as the fraction of poisoned texts among the $N$ poisoned ones that are retrieved for the target question. A higher F1-Score means more poisoned texts are retrieved.
Hyperparameters
We inject 1 adversarial poisoned text for each target question and test on all 100 test data. The results are averaged over 4 random runs.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims we made in the abstract accurately reflect the paper’s contributions and scope, and they are justified in the experiments section.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitation of the paper in the discussion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not have theoretical result in this paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We disclose the details of the algorithm, as well as hyperparameters used to implement our algorithm.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We include the data and code in supplemental material to reproduce our results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify the details in experiment setup section.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report confidence intervals for the experiments we made.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We list the compute resources required.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We confirm to the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss broader impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We don’t think our model has a high risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly credited and are the license and terms of use explicitly mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We document our code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We don’t do crowadsouring and research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We don’t do crowadsouring and research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 Experimentally, we show that our method achieves a high attack success rate on several datasets and using several retrievers, and can generalize to unseen queries and new domains.
-Notably, our method is extremely effective in attacking the ANCE retrieval model, achieving attack success rates that are 17.6% and 13.37% higher on the NQ and MS MARCO datasets, respectively, compared to HotFlip.
+Notably, our method is extremely effective in attacking the ANCE retrieval model, achieving attack success rates that are 15.24% and 17.44% higher on the NQ and MS MARCO datasets, respectively, compared to HotFlip.
 Additionally, we demonstrate AGGD’s potential to replace HotFlip in other adversarial attacks, such as knowledge poisoning of RAG systems.
-[^1] Code can be find in https://github.com/JinyanSu1/AGGD
+Corpus Poisoning via Approximate Greedy Gradient Descent
+Jinyan Su1, Preslav Nakov2, Claire Cardie1 1Department of Computer Science, Cornell University 2Mohamed bin Zayed University of Artificial Intelligence {js3673,ctc9}@cornell.edu, preslav.nakov@mbzuai.ac.ae
 1 Introduction
@@
   • We conduct extensive experiments to show the vulnerability of dense retrievers under AGGD.
-For example, when attacking the ANCE retriever, injecting just one adversarial passage (0.00003% of the NQ dataset corpus and 0.00001% of the MS MARCO dataset corpus) can achieve an attack success rate of 44.35% and 26.16% on these datasets, respectively, improving by 17.6% and 13.37% over HotFlip.
-The generated adversarial passage also possesses the capability to transfer to unseen queries in other domains, with AGGD achieving an attack success rate of 82.28% on the NFCorpus dataset.
+For example, when attacking the ANCE retriever, injecting just one adversarial passage can achieve an attack success rate of 80.92% and 65.68% on these datasets, respectively, improving by 15.24% and 17.44% over HotFlip.
+The generated adversarial passage also possesses the capability to transfer to unseen queries in other domains.
 2 Related Work
@@
 Adversarial Attacks in Retrieval Systems Black-hat search engine optimization, which aims to increase the exposure of certain documents through malicious manipulation, poses a threat by reducing the quality of search results and inundating users with irrelevant pages (Castillo et al., 2011; Liu et al., 2023).
-Previous work has shown that retrieval systems are susceptible to small perturbations: making small edits to a target passage can significantly alter its retrieval rank (Song et al., 2020; Raval & Verma, 2020; Song et al., 2022) for individual or a small set of queries.
+Previous work has shown that retrieval systems are susceptible to small perturbations: making small edits to a target passage can significantly alter its retrieval rank (Song et al., 2020; Raval and Verma, 2020; Song et al., 2022) for individual or a small set of queries.
 More recently, a stronger setting known as corpus poisoning attack has been studied in Zhong et al.
@@
 The retrieval model remains unchanged in a corpus poisoning attack.
-Discrete Optimization Many adversarial attacks in NLP involve discrete optimization, whether in classification tasks (Wallace et al., 2019; Ebrahimi et al., 2018; Song et al., 2021), retrieval systems (Jia & Liang, 2017; Song et al., 2020; Raval & Verma, 2020; Song et al., 2022), or adversarial prompt generation (Zou et al., 2023; Shin et al., 2020; Wen et al., 2024).
+Discrete Optimization Many adversarial attacks in NLP involve discrete optimization, whether in classification tasks (Wallace et al., 2019; Ebrahimi et al., 2018; Song et al., 2021), retrieval systems (Jia and Liang, 2017; Song et al., 2020; Raval and Verma, 2020; Song et al., 2022), or adversarial prompt generation (Zou et al., 2023; Shin et al., 2020; Wen et al., 2024).
 The objective is to find small perturbations to the input to lead the model to make erroneous predictions.
-Due to the discrete nature of texts, directly applying adversarial attack methods from prior computer vision research (Xiao & Wang, 2021; Tolias et al., 2019) is infeasible.
+Due to the discrete nature of texts, directly applying adversarial attack methods from prior computer vision research (Xiao and Wang, 2021; Tolias et al., 2019) is infeasible.
 Instead, many methods build upon HotFlip (Ebrahimi et al., 2018) and approximate the effect of replacing a token using gradients.
@@
 Evaluation Metrics After generating the adversarial passages on the training set and injecting them into the corpus, we evaluate the effectiveness of our attack using the top-$k$ attack success rate (ASR) on test queries.
-Top-$k$ ASR is defined as the percentage of queries for which at least one adversarial passage is retrieved in the top-$k_{r}$ results, i.e., $\text{ASR}=\frac{1}{n_{q}}\sum_{i=1}^{n_{q}}\mathbbm{1}\{a\in\mathcal{R}_{r}(q^{\text{test}}_{i},k_{r})\}$, where $n_{q}$ is the total number of test queries and $\mathcal{R}_{r}(q_{i}^{\text{test}},k_{r})$ is the retriever that returns the top-$k_{r}$ most relevant passages for the test test query $q_{i}^{\text{test}}$.
-$\mathbbm{1}\{a\in r(q_{i}^{\text{test}},k_{r})\}$ is the indicator function, which equals to 1 if $a\in\mathcal{R}(q_{i}^{\text{test}},k_{r})$ and 0 otherwise.
+Top-$k$ ASR is defined as the percentage of queries for which at least one adversarial passage is retrieved in the top-$k_{r}$ results, i.e., $\text{ASR}=\frac{1}{n_{q}}\sum_{i=1}^{n_{q}}\mathbbm{1}\{a\in\mathcal{R}_{r}(q^{\text{test}}_{i},k_{r},\mathcal{C}_{\text{test}})\}$, where $n_{q}$ is the total number of test queries and $\mathcal{R}_{r}(q_{i}^{\text{test}},k_{r},\mathcal{C}_{\text{test}})$ is the retriever that returns the top-$k_{r}$ most relevant passages for the test query $q_{i}^{\text{test}}$.
+$\mathbbm{1}\{a\in\mathcal{R}_{r}(q_{i}^{\text{test}},k_{r},\mathcal{C}_{\text{test}})\}$ is the indicator function, which equals to 1 if $a\in\mathcal{R}_{r}(q_{i}^{\text{test}},k_{r},\mathcal{C}_{\text{test}})$ and 0 otherwise.
 A higher ASR indicates that the model is more vulnerable to attacks, and thus, the attack is more effective.
-We use $k_{r}=1000$ unless specified otherwise.
+We use $k_{r}=20$ to present our result.
+Since ASR depends on the size the the test corpus, to make fair comparisons across different dataset, we randomly sample $|\mathcal{C}_{\text{test}}|=10,000$ passages from the overall corpus pool.
 Additionally, we use retrieval accuracy on the validation data, which is defined as
@@
 Contriever Contriever-MS ANCE DPR-mul DPR-nq
-NQ AGGD 89.57(3.58) 54.53(11.29) 44.35(3.54) 1.47(0.41) 0.14(0.06)
-HotFlip 88.46(0.84) 49.64(6.09) 26.75(1.45) 1.01(0.15) 0.19(0.12)
-Random 74.64(1.78) 18.34(1.48) 7.75(1.88) 0.5(0.21) 0.19(0.1)
-MS MARCO AGGD 75.58(5.81) 13.95(15.16) 26.16(13.3) 2.91(1.01) 2.91(1.93)
-HotFlip 75.0(4.76) 13.95(9.16) 12.79(5.07) 1.74(1.01) 0.58(1.01)
-Random 55.81(6.98) 1.74(3.02) 5.81(3.49) 1.16(2.01) 0.58(1.01)
+NQ AGGD 92.5(2.68) 63.45(8.68) 80.92(4.82) 6.88(1.72) 2.19(0.98)
+Hot Flip 91.08(0.9) 58.43(4.53) 65.68(2.5) 5.4(0.36) 2.03 (0.25)
+Random 80.24(0.92) 32.5(2.61) 31.0(4.3) 3.7(0.93) 1.66 (0.14)
+MS MARCO AGGD 85.47(3.81) 24.42(17.44) 93.6(1.01) 12.79(3.86) 5.23 (1.01)
+Hot Flip 83.72(5.93) 22.67(12.01) 76.16(9.21) 9.88(1.93) 5.81(2.01)
+Random 66.86(7.24) 13.95(13.46) 49.42(11.2) 6.4(2.53) 2.91 (1.93)
 Table 1: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO datasets with 5 retrievers by injecting 1 adversarial passage.
-(Equivalent to increasing the corpus size by 0.00003% for NQ and 0.00001% for MS MARCO.) We highlight the best performing attacking method in bold (Higher ASR indicates better attack performance).
+We highlight the best performing attacking method in bold (Higher ASR indicates better attack performance).
 Results are from 4 random runs with standard deviation in parenthesis.
@@
 All three attack baselines achieve the highest ASR compared to other retrieval models.
-Poisoning with AGGD achieves 89.57% ASR on NQ dataset and 75.58% on MS MARCO dataset, respectively.
-Even using Random perturbation can achieve a relatively high ASR of 74.64% on NQ dataset and 55.81% on MS MARCO dataset.
-(2) Besides achieving comparable results in Contriever and DPR, AGGD is extremely effective in attacking ANCE, improving over the second-best baseline by 17.6% and 13.37% on NQ and MS MARCO, respectively.
+Poisoning with AGGD achieves 92.5% ASR on NQ dataset and 85.47% on MS MARCO dataset, respectively.
+Even using Random perturbation can achieve a relatively high ASR of 80.24% on NQ dataset and 66.86% on MS MARCO dataset.
+(2) Besides achieving comparable results in Contriever and DPR, AGGD is extremely effective in attacking ANCE, improving over the second-best baseline by 15.24% and 17.44% on NQ and MS MARCO, respectively.
 The effectiveness of AGGD in attack ANCE can also be clearly observed during the training.
@@
 (3) Supervised retriever models such as DPR are more challenging to attack.
-Attacking DPR-nq on NQ with only 1 adversarial passage does no better than random perturbation.
-(4) MS MARCO seems to be harder to attack than NQ, likely due to its larger testing corpus, and thus, lower poison rate.
+Attacking DPR-nq on NQ with only 1 adversarial passage are just slightly better than random perturbation.
 Target Domain Source Domain Methods Retriever
 Contriever Contriever-MS ANCE DPR-mul DPR-nq
-FiQA-2018 NQ AGGD 62.04(17.22) 3.78(1.39) 4.09(1.18) - -
-HotFlip 65.28(9.06) 2.08(0.55) 2.89(0.62) 0.23(0.23) 0.15(0.11)
-Random 37.38(2.76) 0.27(0.13) 0.96(0.33) 0.15(0.11) 0.31(0.11)
-MS MARCO AGGD 84.57(1.97) 14.16(21.18) 11.46(1.61) 1.89(0.56) 1.97(1.24)
-HotFlip 75.5(7.87) 3.24(2.06) 7.02(0.56) 1.2(0.6) 1.81(1.18)
-Random 58.76(6.12) 0.23(0.32) 3.51(1.06) 0.46(0.24) 0.42(0.13)
-NFCorpus NQ AGGD 51.32(7.79) 20.98(5.59) 49.54(6.21) 12.62(2.09) 16.25(0.56)
-HotFlip 45.9(11.47) 13.24(2.6) 46.13(2.22) 13.85(3.57) 17.41(1.63)
-Random 28.79(6.24) 8.67(2.19) 27.01(5.86) 10.29(2.03) 16.18(0.46)
-MS MARCO AGGD 56.58(14.85) 22.91(18.3) 82.28(3.41) 38.16(5.66) 40.63(4.49)
-HotFlip 54.49(6.28) 17.96(10.63) 65.56(4.99) 29.18(3.16) 34.98(8.55)
-Random 32.97(5.71) 8.59(8.34) 55.73(7.11) 23.61(4.15) 23.14(2.81)
-Quora NQ AGGD 57.36(14.93) 11.24(4.69) 24.48(3.48) 0.36(0.06) 0.27(0.14)
-HotFlip 61.82(4.97) 8.09(4.42) 20.43(0.73) 0.3(0.14) 0.2(0.06)
-Random 30.09(4.01) 1.03(0.1) 10.88(0.94) 0.17(0.02) 0.27(0.05)
-MS MARCO AGGD 73.42(4.37) 8.25(12.74) 37.07(2.52) 3.32(1.32) 3.37(1.07)
-HotFlip 69.62(8.69) 2.75(1.59) 27.92(2.31) 2.03(0.95) 2.31(0.34)
-Random 42.14(7.59) 0.37(0.45) 18.39(0.85) 0.8(0.2) 0.7(0.54)
-SCIDOCS NQ AGGD 21.18(7.91) 11.75(7.84) 9.55(1.72) 0.87(0.26) 0.4(0.1)
-HotFlip 27.15(13.45) 14.1(4.41) 7.18(1.0) 1.18(0.4) 0.35(0.11)
-Random 12.82(3.62) 0.73(0.58) 2.9(0.8) 0.57(0.37) 0.27(0.08)
-MS MARCO AGGD 23.25(6.88) 14.22(20.72) 31.32(7.33) 3.92(1.7) 1.13(0.45)
-HotFlip 24.98(6.23) 11.98(7.4) 21.25(4.37) 3.05(1.34) 0.85(0.32)
-Random 15.02(1.6) 1.23(1.74) 10.58(5.27) 2.35(1.27) 0.23(0.08)
-SciFact NQ AGGD 25.83(4.19) 1.58(0.86) 10.83(2.88) 0.17(0.17) -
-HotFlip 28.83(14.18) 1.17(0.83) 6.75(0.92) 0.42(0.43) -
-Random 9.58(4.44) - 1.42(0.86) - -
-MS MARCO AGGD 31.83(10.62) 3.75(5.01) 48.0(7.04) 3.5(0.9) 1.58(0.49)
-HotFlip 25.0(7.06) 2.5(1.89) 28.17(8.07) 2.83(0.5) 1.42(1.44)
-Random 12.08(6.19) 0.17(0.17) 13.42(7.98) 1.75(1.09) 0.67(0.71)
-Table 2: Out-of-domain top-50 attack success rate with only 1 adversarial passage.
+FiQA-2018 NQ AGGD 64.93(15.9) 4.94(1.68) 6.33(1.91) 0.15(0.11) 0.12(0.13)
+Hot Flip 69.68(8.66) 3.05(0.6) 4.13(0.77) 0.46(0.55) 0.62(0.15)
+Random 41.05(2.73) 0.35(0.13) 1.66(0.49) 0.27(0.17) 0.46(0.19)
+MS MARCO AGGD 86.54(1.88) 15.66(22.54) 17.98(1.44) 3.59(1.34) 3.36(1.79)
+Hot Flip 78.51(6.96) 4.59(2.89) 11.5(0.79) 2.16(1.13) 3.2(1.72)
+Random 62.81(6.07) 0.42(0.57) 5.48(1.36) 0.73(0.2) 0.93(0.24)
+NFCorpus NQ AGGD 46.83(8.1) 17.1(4.3) 37.69(4.2) 8.51(1.29) 12.23(0.46)
+Hot Flip 42.34(10.96) 10.91(2.77) 34.68(1.71) 10.22(2.62) 12.85(0.99)
+Random 24.38(5.02) 6.58(1.43) 18.34(4.09) 7.43(1.3) 11.3(0.56)
+MS MARCO AGGD 52.55(13.94) 18.73(16.3) 69.81(3.57) 26.16(4.51) 28.17(4.02)
+Hot Flip 49.92(5.26) 14.24(8.46) 52.55(6.86) 19.27(2.31) 23.06(7.02)
+Random 30.03(6.42) 6.42(6.45) 42.26(6.23) 14.47(3.59) 14.78(1.06)
+Quora NQ AGGD 73.92(12.16) 25.1(7.03) 77.72(5.57) 4.85(1.13) 6.39(2.1)
+Hot Flip 78.99(3.56) 18.92(5.84) 72.37(2.13) 3.36(1.02) 7.2(1.45)
+Random 49.44(4.18) 6.12(0.63) 53.09(4.89) 2.52(0.47) 6.66(1.51)
+MS MARCO AGGD 86.92(2.49) 18.83(21.47) 91.11(1.48) 22.29(4.24) 25.96(5.05)
+Hot Flip 83.26(5.87) 12.82(7.64) 84.18(2.25) 15.38(4.14) 23.07(3.67)
+Random 62.57(6.63) 3.81(4.12) 74.06(0.75) 8.82(1.14) 12.29(3.33)
+SCIDOCS NQ AGGD 21.05(7.78) 11.95(7.99) 9.78(1.97) 0.88(0.29) 0.38(0.08)
+Hot Flip 27.02(13.4) 14.32(4.48) 7.1(0.87) 1.18(0.4) 0.35(0.11)
+Random 12.88(3.76) 0.78(0.61) 3.05(0.93) 0.6(0.37) 0.3(0.07)
+MS MARCO AGGD 23.22(6.99) 14.3(20.69) 31.9(7.2) 3.88(1.62) 1.08(0.42)
+Hot Flip 24.98(6.33) 12.25(7.52) 21.88(4.33) 3.12(1.4) 0.85(0.32)
+Random 15.0(1.53) 1.27(1.83) 10.62(5.11) 2.48(1.38) 0.2(0.07)
+SciFact NQ AGGD 22.92(4.46) 1.0(0.62) 6.75(1.92) - -
+Hot Flip 24.5(12.16) 0.92(0.72) 3.58(0.98) 0.17(0.29) -
+Random 8.25(3.42) - 0.5(0.37) - -
+MS MARCO AGGD 29.08(10.43) 2.67(4.43) 30.33(6.47) 2.42(0.6) 0.5(0.29)
+Hot Flip 21.58(6.52) 1.25(1.11) 16.5(4.78) 2.0(0.24) 0.67(0.71)
+Random 9.83(5.46) - 7.33(4.96) 1.17(0.96) 0.17(0.17)
+Table 2: Out-of-domain top-20 attack success rate with only 1 adversarial passage.
 Results are averaged over 4 random runs with standard deviation shown in parenthesis.
@@
 In Table 2, we use adversarial passages generated from training set of NQ and MS MARCO and insert them into the corpora of retrieval tasks in other domains.
-Since the out-of-domain datasets have smaller corpora compared to NQ and MS MARCO, we use top $k_{r}=50$ ASR (rather than $k_{r}=1000$ used for in-domain attacks) to measure the effectiveness of the attack method.
 We found that (1) Compared to adversarial passages generated from NQ dataset, those trained from MS MARCO generally perform better in out-of-domain attacks, possibly because MS MARCO contains more training data.
 (2) Contriever models are still the most vulnerable to corpus poisoning attacks.
-For example, inserting a single adversarial passage generated by AGGD into FiQA-2018 achieves a top-$k_{r}=50$ ASR of 84.57%, and inserting it into Quora can trick the model into returning the adversarial passage in the top-50 retrieved passage for 73.42% of the queries.
-(3) NFCorpus is an easier domain to attack.
-AGGD achieves over 38% ASR even in DPR-mul and DPR-nq, which is surprising since, as shown in the in-domain attack results in Table 1, attacking DPR-mul and DPR-nq is extremely hard , even with adversarial passage trained from in-domain data.
+For example, inserting a single adversarial passage generated by AGGD into FiQA-2018 achieves a top-$k_{r}=20$ ASR of 86.54%, and inserting it into Quora can trick the model into returning the adversarial passage in the top-20 retrieved passage for 86.92% of the queries.
+(3) Quora are easier domains to attack.
+AGGD achieves over 20% ASR even in DPR-mul and DPR-nq, which is surprising since, as shown in the in-domain attack results in Table 1, attacking DPR-mul and DPR-nq is extremely hard , even with adversarial passage trained from in-domain data.
+(Though NFCorpus has high attack success rate, NFCorpus also has a smaller testing corpus).
 5.3 Analysis and Ablation Study
@@
 The effect of the candidate set size would be more pronounced when a larger range of candidate set sizes is considered.
+LLaMa-2-7B LLaMa-2-13B Vicuna-7B Vicuna-13B Vicuna-33B GPT-3.5 GPT-4
+MS MARCO
+AGGD 0.81(0.00) 0.79(0.01) 0.73(0.01) 0.72(0.01) 0.67(0.01) 0.80(0.01) 0.88(0.01)
+HotFlip 0.80(0.02) 0.77(0.02) 0.69(0.02) 0.73(0.04) 0.69(0.02) 0.79(0.01) 0.86(0.01)
+NQ
+AGGD 0.85(0.00) 0.94(0.00) 0.81(0.00) 0.82(0.00) 0.68(0.00) 0.92(0.00) 0.97(0.00)
+HotFlip 0.83(0.00) 0.95(0.00) 0.80(0.00) 0.81(0.00) 0.67(0.00) 0.87(0.01) 0.96(0.01)
+Table 3: Comparing Attack success rate of PoisonedRAG (white-box) using HotFlip and AGGD to find the adversarial texts.
 The effect of the token length $m$
@@
 We find that AGGD achieves a comparable or better attack success rate than Hot Flip in both MS MARCO and NQ dataset on multiple LLMs.
-LLaMa-2-7B LLaMa-2-13B Vicuna-7B Vicuna-13B Vicuna-33B GPT-3.5 GPT-4
-MS MARCO
-AGGD 0.81(0.00) 0.79(0.01) 0.73(0.01) 0.72(0.01) 0.67(0.01) 0.80(0.01) 0.88(0.01)
-HotFlip 0.80(0.02) 0.77(0.02) 0.69(0.02) 0.73(0.04) 0.69(0.02) 0.79(0.01) 0.86(0.01)
-NQ
-AGGD 0.85(0.00) 0.94(0.00) 0.81(0.00) 0.82(0.00) 0.68(0.00) 0.92(0.00) 0.97(0.00)
-HotFlip 0.83(0.00) 0.95(0.00) 0.80(0.00) 0.81(0.00) 0.67(0.00) 0.87(0.01) 0.96(0.01)
-Table 3: Comparing Attack success rate of PoisonedRAG (white-box) using HotFlip and AGGD to find the adversarial texts.
 6 Conclusion
@@
 Additional experiments on other adversarial attacks indicate the potential of AGGD as a competitive alternative to the widely used HotFlip.
-7 Limitations and Broader Impact
-Though we use corpus poisoning attack as the main example to deliver our attack algorithm AGGD, our proposed attack can be used for various attack scenarios, thus, we didn’t specifically investigate defenses of AGGD for corpus poison attack.
-Since our method assembles to HotFlip, existing defenses to these attacks should be able to defend our attack to some extend.
-However, as our attack is stronger than HotFlip, it remains to understand how well these defense works.
+7 Limitations and Future Work
+While we use corpus poisoning attacks to showcase our AGGD algorithm, the proposed attack framework is versatile and applicable to a wide range of adversarial attack scenarios, as many of these can be formulated as discrete optimization problems.
+However, the generated output sequences often lack semantic coherence, making the adversarial corpus easily detectable and filterable.
+A promising direction for future work is to reformulate the problem as a constrained optimization problem, focusing on producing semantically meaningful adversarial passages that are more difficult to defend against, even if this may compromise the overall attack success rate.
+8 Ethics Statement
+Our work studies the vulnerability of dense retriever and corpus poison attack.
+The propose attack AGGD shows higher attach success rate especially for ANCE model, compared to previous HotFlip attack, which could be used for spread garbage information.
+Future research based on this paper should be exercised with caution and consider the potential consequences.
 Appendix A Additional Experiments
@@
 In-Domain Attack with 10 Adversarial Passages
-Table 4 shows the in-domain attack results of inserting 10 adversarial passages into the NQ and MS MARCO datasets, with a poison rate of around $0.0003\%$ of the NQ corpus and $0.0001\%$ of the MS MARCO corpus.
-Similar to injecting only 1 adversarial passage, we found that (1) The pretrained Contriever model is easy to attack: even with just 10 adversarial passages, all three baselines (AGGD, HotFlip and Random perturbation) successfully attack more than 90% of the queries in both NQ and MS MARCO datasets, tricking the Contriever into returning the adversarial passage among the top-1000 retrieved results.
-For NQ dataset, even random perturbation achieves a high ASR of 97.75%, while the other two methods using gradient information achieves higher ASR of $>99\%$.
-(2) AGGD still outperforms HotFlip on ANCE, with improvements of 11.95% and 25.58% on NQ and MS MARCO datasets, respectively.
-(3) Though attacking DPR-nq on NQ with only 1 adversarial passage does no better than random perturbation (as shown in Table 1), increasing the number of adversarial passage to 10 allows both AGGD and HotFlip to perform better than random perturbation, with ASR improvement of 10.08% and 6.67% over the Random perturbation baseline.
+Table 4 shows the in-domain attack results of inserting 10 adversarial passages into the NQ and MS MARCO datasets.
+Similar to injecting only 1 adversarial passage, we found that (1) The pretrained Contriever model is easy to attack: even with just 10 adversarial passages, all three baselines (AGGD, HotFlip and Random perturbation) successfully attack more than 80% of the queries in both NQ and MS MARCO datasets, tricking the Contriever into returning the adversarial passage among the top-20 retrieved results.
+For NQ dataset, even random perturbation achieves a high ASR of 80%, while the other two methods using gradient information achieves higher ASR of $>91\%$.
+(2) AGGD still outperforms HotFlip on ANCE, with improvements of 18.62% and 25.58% on NQ and MS MARCO datasets, respectively.
 Dataset Methods Retriever
 Contriever Contriever-MS ANCE DPR-mul DPR-nq
-NQ AGGD 99.46(0.01) 86.76(1.48) 85.53(0.42) 23.41(0.03) 16.87(0.07)
-HotFlip 99.06(0.07) 86.72(2.25) 73.58(1.97) 25.77(1.32) 13.46(0.42)
-Random 97.75(0.04) 78.06(2.56) 55.87(1.35) 15.34(0.74) 6.79(0.07)
-MS MARCO AGGD 96.51(1.16) 54.65(1.16) 72.09(0.0) 13.95(0.0) 5.81(3.49)
-HotFlip 96.51(3.49) 59.3(3.49) 46.51(2.33) 11.63(0.0) 2.33(0.0)
-Random 91.86(1.16) 27.91(2.33) 30.23(9.3) 5.81(3.49) 1.16(1.16)
-Table 4: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO dataset with 5 retrievers by injecting 10 adversarial passages.
-(adding 0.0003% corpus to NQ and 0.0001% to MS MARCO dataset).
+NQ AGGD 91.57(3.5) 69.42(1.38) 82.08(2.01) 8.23(1.48) 1.9 (0.07)
+Hot Flip 91.6(0.87) 57.37(5.95) 63.46(1.55) 5.08(0.1) 2.19(0.25)
+Random 80.49(0.83) 33.73(2.27) 27.3(2.74) 3.94(1.22) 1.65 (0.06)
+MS MARCO AGGD 83.72(4.65) 13.95(13.95) 93.02(0.0) 11.63(2.33) 5.81 (1.16)
+Hot Flip 88.37(2.33) 13.95(11.63) 67.44(2.33) 10.46(1.16) 6.98(2.33)
+Random 65.12(9.3) 1.16(1.16) 52.33(15.12) 8.14(1.16) 3.49 (1.16)
+Table 4: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO dataset with 5 retrievers by injecting 10 adversarial passages).
 The best-performing attacking method is highlighted with in bold (higher ASR indicates better attack performance).
@@
 Contriever Contriever-MS ANCE DPR-mul DPR-nq
-FiQA-2018 NQ AGGD 90.97(0.08) 25.08(7.02) 12.27(0.39) 4.32(0.0) 2.7(0.85)
-HotFlip 79.55(1.77) 20.14(2.86) 8.33(1.7) 1.77(1.31) 1.62(0.54)
-Random 79.17(0.0) 14.43(4.55) 5.56(0.31) 1.23(0.0) 1.23(0.15)
-MS MARCO AGGD 97.38(0.31) 45.99(25.31) 25.69(1.16) 21.22(7.95) 14.35(6.33)
-HotFlip 94.06(1.47) 50.08(3.32) 18.75(0.39) 13.81(5.17) 20.52(0.77)
-Random 83.41(0.23) 5.25(0.31) 8.33(0.93) 10.57(5.32) 4.71(3.01)
-NFCorpus NQ AGGD 83.13(2.01) 43.5(2.32) 75.54(3.72) 25.23(6.97) 25.23(2.94)
-HotFlip 88.85(0.62) 42.88(3.56) 63.0(1.39) 31.89(1.24) 27.09(0.16)
-Random 68.27(0.46) 29.72(4.02) 48.3(2.48) 22.29(0.93) 21.21(2.01)
-MS MARCO AGGD 97.06(0.15) 62.07(25.85) 95.2(0.46) 72.29(7.9) 64.24(4.49)
-HotFlip 90.56(4.8) 78.02(9.6) 89.78(0.93) 69.35(11.15) 74.77(5.42)
-Random 88.85(3.1) 63.78(1.86) 81.58(1.7) 47.21(3.87) 61.92(8.98)
-Quora NQ AGGD 85.62(0.13) 35.84(2.36) 42.12(0.82) 2.7(0.32) 3.53(0.89)
-HotFlip 77.47(1.67) 32.46(9.31) 33.01(2.55) 3.84(0.03) 4.04(0.13)
-Random 67.66(5.7) 23.15(5.66) 26.36(0.42) 1.64(0.09) 2.36(0.47)
-MS MARCO AGGD 91.8(2.98) 34.08(5.24) 48.4(0.52) 12.75(0.04) 12.0(1.43)
-HotFlip 92.31(3.08) 40.83(3.2) 41.58(0.04) 12.93(0.48) 13.66(0.03)
-Random 65.28(3.2) 11.79(3.52) 30.69(1.66) 6.43(0.13) 7.09(0.95)
-SCIDOCS NQ AGGD 53.45(7.45) 45.7(17.7) 33.0(0.9) 1.75(0.35) 0.45(0.05)
-HotFlip 49.7(2.2) 40.95(5.15) 25.4(5.7) 3.2(0.1) 1.0(0.2)
-Random 37.7(4.1) 17.95(1.35) 13.25(0.95) 2.2(0.5) 0.55(0.25)
-MS MARCO AGGD 65.4(2.5) 41.0(10.1) 45.25(2.15) 24.5(2.2) 6.4(0.0)
-HotFlip 51.35(4.65) 67.55(8.25) 36.1(0.2) 21.85(2.95) 7.95(0.75)
-Random 39.05(4.55) 13.6(4.1) 20.25(0.65) 14.7(2.9) 2.5(1.1)
-SciFact NQ AGGD 66.83(5.5) 16.33(9.67) 56.83(9.5) 1.83(1.5) 0.33(0.33)
-HotFlip 63.5(12.17) 15.67(3.0) 27.0(8.0) 2.83(1.17) 1.0(0.0)
-Random 41.17(4.5) 5.0(3.0) 12.67(0.67) 1.0(0.0) 0.67(0.0)
-MS MARCO AGGD 93.83(2.5) 49.83(13.17) 72.5(1.17) 17.33(5.67) 17.5(10.5)
-HotFlip 89.5(3.17) 33.5(8.83) 58.83(6.17) 24.5(1.17) 21.17(10.83)
-Random 80.83(8.17) 13.67(4.0) 32.67(8.0) 8.33(2.33) 14.5(2.17)
-Table 5: Out-of-domain top-50 attack success rate with 10 adversarial passages.
+FiQA-2018 NQ AGGD 92.67(0.23) 30.86(7.87) 19.52(1.47) 6.02(0.15) 4.71(1.47)
+Hot Flip 84.1(0.77) 24.23(2.31) 13.35(2.7) 3.94(2.55) 3.24(1.23)
+Random 83.1(0.39) 17.82(5.32) 8.26(0.54) 2.7(0.23) 2.31(0.46)
+MS MARCO AGGD 97.69(0.31) 50.54(25.23) 37.19(0.31) 27.31(7.87) 20.6(7.64)
+Hot Flip 95.22(1.08) 56.48(3.4) 28.16(2.08) 19.37(6.4) 28.47(1.31)
+Random 86.73(0.15) 6.94(0.15) 13.27(1.39) 15.66(7.48) 7.25(4.63)
+NFCorpus NQ AGGD 78.02(1.24) 35.76(3.25) 64.55(3.87) 18.11(4.8) 18.89(3.1)
+Hot Flip 85.45(0.31) 37.0(3.87) 48.76(1.08) 20.43(0.31) 20.59(1.39)
+Random 63.78(1.86) 24.46(3.41) 35.6(0.62) 16.72(1.55) 15.79(1.55)
+MS MARCO AGGD 95.36(0.62) 57.9(26.01) 88.39(0.15) 57.89(8.05) 52.01(3.72)
+Hot Flip 87.62(5.88) 70.28(12.69) 80.34(1.39) 59.75(10.84) 66.72(4.8)
+Random 85.45(3.1) 51.39(5.57) 65.17(2.63) 34.98(1.24) 52.48(8.2)
+Quora NQ AGGD 96.18(0.03) 63.14(3.93) 91.52(0.89) 20.96(3.63) 25.8(5.81)
+Hot Flip 92.22(1.08) 57.88(10.45) 84.9(2.27) 24.82(0.15) 32.19(1.7)
+Random 87.85(2.18) 46.28(7.64) 81.72(0.25) 15.84(0.97) 22.28(3.28)
+MS MARCO AGGD 97.87(0.93) 61.72(7.29) 94.52(0.17) 45.92(1.0) 51.5(3.35)
+Hot Flip 98.13(0.79) 69.88(3.63) 91.5(0.16) 43.86(2.64) 51.6(3.73)
+Random 83.97(3.6) 35.24(4.43) 85.48(2.07) 32.88(0.31) 38.36(6.64)
+SCIDOCS NQ AGGD 53.75(7.05) 46.4(17.7) 33.75(1.25) 1.7(0.3) 0.45(0.05)
+Hot Flip 49.75(2.45) 40.7(5.4) 26.45(6.05) 3.35(0.15) 0.95(0.15)
+Random 37.8(4.3) 18.25(1.75) 13.3(0.9) 2.3(0.4) 0.6(0.2)
+MS MARCO AGGD 65.5(2.6) 41.65(9.85) 46.2(1.9) 24.9(2.2) 6.7(0.2)
+Hot Flip 51.45(4.75) 68.1(8.3) 36.95(0.35) 22.4(2.9) 8.3(1.1)
+Random 39.2(4.5) 13.9(4.3) 20.9(0.6) 15.55(3.35) 2.45(1.05)
+SciFact NQ AGGD 60.33(6.0) 12.5(7.17) 39.33(9.33) 0.5(0.5) 0.17(0.17)
+Hot Flip 58.83(12.83) 11.83(2.17) 16.5(3.5) 1.67(0.67) 0.5(0.17)
+Random 34.33(4.33) 3.17(2.17) 8.17(1.5) 0.33(0.0) 0.5(0.17)
+MS MARCO AGGD 91.5(4.5) 41.5(8.5) 58.17(0.83) 12.0(5.0) 12.0(7.67)
+Hot Flip 85.83(5.17) 29.17(9.5) 42.33(4.0) 15.17(0.5) 14.67(7.33)
+Random 75.5(9.5) 8.0(1.33) 21.5(6.5) 5.67(1.67) 10.33(1.0)
+Table 5: Out-of-domain top-20 attack success rate with 10 adversarial passages.
 Due to the computational constraints, results are averaged over 2 random runs, with standard deviations shown in parentheses.
 The combinations of attack and source dataset that achieve the highest ASR for each target domain and retriever are highlighted in bold.
+B.2 In-domain attack success rate with different $k$.
+Dataset Method Retriever
+Contriever Contriever-MS ANCE DPR-mul DPR-nq
+Top-1 NQ AGGD 85.06(5.06) 44.84(12.65) 16.36(1.52) 0.39(0.15) -
+Hot Flip 83.77(1.43) 39.74(7.81) 8.66(0.62) 0.21(0.08) 0.03 (0.02)
+Random 66.64(2.61) 8.82(0.87) 1.89(0.5) 0.07(0.04) 0.04(0.03)
+MS MARCO AGGD 69.19(6.65) 11.05(12.67) 24.42(7.45) 1.16(1.16) 2.91(1.93)
+Hot Flip 71.51(4.15) 12.79(8.47) 10.47(2.01) 2.33(1.64) 0.58 (1.01)
+Random 52.33(6.26) 1.74(1.93) 5.23(4.47) - 0.58 (1.01)
+Top-5 NQ AGGD 89.84(3.53) 55.63(10.96) 49.56(3.78) 1.9(0.55) 0.27 (0.16)
+Hot Flip 88.69(0.72) 50.64(5.96) 31.36(1.16) 1.29(0.26) 0.28(0.1)
+Random 75.28(1.8) 20.07(1.44) 10.12(1.82) 0.71(0.28) 0.24 (0.11)
+MS MARCO AGGD 81.98(5.03) 20.35(16.72) 66.28(5.33) 8.72(3.02) 3.49(1.16)
+Hot Flip 80.23(6.04) 19.77(10.2) 36.05(9.08) 3.49(1.16) 2.33 (0.0)
+Random 62.21(7.6) 8.14(7.07) 20.93(4.03) 3.49(2.01) 0.58 (1.01)
+Top-10 NQ AGGD 91.28(3.03) 59.39(9.83) 66.43(4.42) 3.62(1.02) 0.82(0.38)
+Hot Flip 89.85(0.82) 54.52(5.41) 47.55(1.73) 2.69(0.24) 0.72 (0.2)
+Random 77.8(1.45) 25.91(2.07) 18.8(3.26) 1.61(0.45) 0.66 (0.1)
+MS MARCO AGGD 84.88(3.86) 22.67(17.35) 85.47(2.53) 9.3(3.68) 4.07(1.01)
+Hot Flip 81.4(5.45) 20.35(10.58) 57.56(9.36) 4.65(1.64) 2.91 (1.01)
+Random 65.12(7.89) 11.63(10.78) 35.47(8.28) 4.65(2.85) 1.16 (1.16)
+Top-20 NQ AGGD 92.5(2.68) 63.45(8.68) 80.92(4.82) 6.88(1.72) 2.19(0.98)
+Hot Flip 91.08(0.9) 58.43(4.53) 65.68(2.5) 5.4(0.36) 2.03 (0.25)
+Random 80.24(0.92) 32.5(2.61) 31.0(4.3) 3.7(0.93) 1.66 (0.14)
+MS MARCO AGGD 85.47(3.81) 24.42(17.44) 93.6(1.01) 12.79(3.86) 5.23 (1.01)
+Hot Flip 83.72(5.93) 22.67(12.01) 76.16(9.21) 9.88(1.93) 5.81(2.01)
+Random 66.86(7.24) 13.95(13.46) 49.42(11.2) 6.4(2.53) 2.91 (1.93)
+Top-50 NQ AGGD 93.87(2.26) 68.69(6.66) 93.7(3.21) 15.59(3.41) 7.31(2.02)
+Hot Flip 92.56(0.82) 63.28(3.64) 85.93(2.4) 13.37(1.1) 6.84 (1.1)
+Random 83.01(0.46) 42.35(3.55) 53.01(5.31) 9.6(1.76) 5.46 (0.51)
+MS MARCO AGGD 87.21(5.33) 31.98(20.3) 99.42(1.01) 22.67(5.55) 13.37(3.02)
+Hot Flip 84.88(4.19) 32.56(18.16) 95.35(1.64) 19.77(3.49) 10.46 (5.07)
+Random 69.19(4.47) 19.19(17.35) 75.58(5.33) 12.79(3.49) 4.65 (1.64)
+Top-100 NQ AGGD 94.77(2.08) 72.59(5.28) 98.2(1.31) 27.94(5.7) 18.0(3.6)
+Hot Flip 93.58(0.66) 67.27(2.97) 94.94(1.63) 24.48(2.26) 16.8 (1.79)
+Random 85.11(0.41) 50.86(4.34) 71.57(5.81) 18.6(3.15) 12.81 (1.34)
+MS MARCO AGGD 89.54(4.79) 40.12(24.36) 100.0(0.0) 41.28(5.03) 20.93(4.35)
+Hot Flip 86.05(4.93) 40.12(18.12) 100.0(0.0) 31.98(5.04) 18.02 (5.78)
+Random 70.93(2.6) 23.26(18.09) 90.12(6.01) 18.02(2.53) 9.88 (4.47)
+Table 6: In-domain attack success rate (ASR) of AGGD on NQ and MS MARCO datasets with 5 retrievers by injecting 1 adversarial passage with varying $k=\{1,5,10,20,50,100\}$.
+Results are from 4 random runs with standard deviation in parenthesis.For the ease of presentation, we omit the results through ‘-’ if top-$k$ ASR is smaller than 0.1%.
+B.3 Transferability over different retrievers.
+Source Retriever Method Target Retriever
+Contriever Contriever-MS ANCE DPR-mul DPR-nq
+contriever AGGD 97.14 26.51 - - 1.2
+Hot Flip 96.05 20.9 - 0.18 0.48
+Random 90.5 11.8 - 0.14 1.39
+contriever-msmarco AGGD 27.76 83.71 - - 0.28
+Hot Flip 25.23 79.05 - 1.02 2.28
+Random 26.24 78.38 - 2.05 7.16
+ance AGGD 26.43 4.88 100.0 - -
+Hot Flip 28.48 4.92 $9$
+Random 9.86 6.39 99.3 - -
+dpr-multi AGGD 15.78 3.67 - 87.84 22.47
+Hot Flip 23.43 7.7 - 84.18 38.41
+Random 18.43 4.15 - 79.35 10.41
+dpr-single AGGD 10.85 9.99 - 29.16 89.46
+Hot Flip 11.87 9.45 - 27.91 87.72
+Random 22.02 7.37 - 20.5 82.55
+Table 7: Attack transferability across models on NQ dataset.
 Appendix C Experimental Detail
@@
 (2017), WebQuestions Berant et al.
-(2013) and CuratedTREC Baudiš & Šedivỳ (2015)) and the single NQ model (DPR-nq).
+(2013) and CuratedTREC Baudiš and Šedivỳ (2015)) and the single NQ model (DPR-nq).
   • ANCE Xiong et al.
@@
 We inject 1 adversarial poisoned text for each target question and test on all 100 test data. The results are averaged over 4 random runs.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The claims we made in the abstract accurately reflect the paper’s contributions and scope, and they are justified in the experiments section.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We discuss the limitation of the paper in the discussion section.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: We do not have theoretical result in this paper.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We disclose the details of the algorithm, as well as hyperparameters used to implement our algorithm.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We include the data and code in supplemental material to reproduce our results.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: We specify the details in experiment setup section.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: We report confidence intervals for the experiments we made.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We list the compute resources required.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We confirm to the code of ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: We discuss broader impacts.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: We don’t think our model has a high risk of misuse.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We properly credited and are the license and terms of use explicitly mentioned and properly respected.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We document our code.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: We don’t do crowadsouring and research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We don’t do crowadsouring and research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+Appendix E Examples of the output passage from AGGD
+Examples from attacking Contriever Examples from attacking contriever-MS
+["##nius", "half", "melting", "definite", "[MASK]", "favor", "gma", "who", "should", "jew", "rain", "##pi", "##ntial", "upper", "kevin", "perrin", "##gren", "blew", "demonstrators", "remnants", "shelters", "that", "##ntial", "hue", "lest", "rainfall", "where", "rains", "rain", "heavy"] ["hostage", "tightly", "handful", "herbs", "where", "packets", "were", "##ssen", "dealers", "overnight", "symbol", "##atic", "##ised", "adventure", "bail", "##alis", "##kari", "[CLS]", "convicted", "ga", "##oit", "peace", "restore", "lifespan", "discrimination", "(", "maha", "##ter", "bank", "##ees"]
+Table 10: Examples of adversarial passages